1. For purposes of this problem, let us define a "checkerboard matrix" to be a square matrix A = [aij] such that

a Subscript i j Baseline equals Start Layout left-brace 1st Row 1st Column 1 if i+j is even 2nd Row 1st Column 0 if i+j is odd End Layout

Find the rank and nullity of the following checkerboard matrix.

The 6 × 6 checkerboard matrix.

a. Rank (A) =
b. nullity (A) =

2. For purposes of this problem, let us define a "checkerboard matrix" to be a square matrix Upper A equals left-bracket a Subscript i j Baseline right-bracket such that

a Subscript i j Baseline equals Star tLayout left-brace 1st Row 1st Column -2 if i plus j is even 2nd Row 1st Column 0 if i plus j is odd End Layout

Find the rank and nullity of the n times n checkerboard matrix for n greater-than-or-equal-to 2.

a. Rank (A) =
b. nullity (A) =
1
Expert's answer
2020-01-16T11:30:09-0500

1. "A=\begin{bmatrix}
 1&0&1&0&1&0\\
 0&1&0&1&0&1\\
1&0&1&0&1&0\\
0&1&0&1&0&1\\
1&0&1&0&1&0\\
0&1&0&1&0&1
\end{bmatrix}" is the 6x6 checkerboards matrix.

Applying "R_6 \gets R_6-R_2" ; "R_4 \gets R_4-R_2" ; "R_5 \gets R_5-R_1" and "R_3 \gets R_3-R_1" ; we get;

"A=\begin{bmatrix}
 1&0&1&0&1&0\\
 0&1&0&1&0&1\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0
\end{bmatrix}" . Clearly, rank(A)=2.

Nullity is given by; Ax=0; where x represents Null Space of A.

"\begin{bmatrix}
 1&0&1&0&1&0\\
 0&1&0&1&0&1\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0
\end{bmatrix}\begin{bmatrix}
 x_1 \\x_2\\x_3\\x_4\\x_5\\x_6
\end{bmatrix}
=\begin{bmatrix}
 0 \\0\\0\\0\\0\\0
\end{bmatrix}"

"\implies x_1+x_3+x_5=0;\implies x_1=-x_3-x_5;"

"\implies x_2+x_4+x_6=0;\implies x_2=-x_4-x_6"

"\begin{bmatrix}
 x_1 \\x_2\\x_3\\x_4\\x_5\\x_6
\end{bmatrix}=\begin{bmatrix}
 -x_3-x_5\\-x_4-x_6\\x_3\\x_4\\x_5\\x_6
\end{bmatrix}" . Thus, 4 independent parameters, "x_3,x_4,x_5,x_6\implies Nullity(A)=4"



2."A=\begin{bmatrix}
 -2&0&-2&0&-2&0\\
 0&-2&0&-2&0&-2\\
-2&0&-2&0&-2&0\\
0&-2&0&-2&0&-2\\
-2&0&-2&0&-2&0\\
0&-2&0&-2&0&-2
\end{bmatrix}" is the checkerboard matrix (for n=6).

Applying "R_6 \gets R_6-R_2; R_4 \gets R_4-R_2; R_5 \gets R_5-R_1and R_3 \gets R_3-R_1" we get;

"A=\begin{bmatrix}
 -2&0&-2&0&-2&0\\
 0&-2&0&-2&0&-2\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0\\
0&0&0&0&0&0
\end{bmatrix}" Clearly, rank(A)=2. Moreover this applies to all matrices A for n>2. As we can perform the same row transformations again and again.


"\implies Rank(A)=2"

Similarly;"Nullity(A)=4"